# A Framework for Guided Motion Planning

### v1

@@ Abstract
I Introduction
A general paradigm in Artificial Intelligence is that one can solve provably worst case intractable search problems using heuristics that are effective in the average case, i.e., in most real world problems. This same principle holds true in robotics, where much of the difficulty stems from the geometry and the implicit nature of the search space, called the configuration space (c-space). The family of sampling based motion planning (SBMP) algorithms searches the continuous c-space using random sampling and local extensions in order to find collision free paths.
Most such algorithms implicitly perform guided search. Early algorithms, such as Rapidly-exploring Random Trees (RRT) [23], showed that exploration through Voronoi bias, i.e., choosing which node to expand with probability proportional to the volume of its Voronoi cell, is highly effective in the absence of arbitrarily narrow passages. As such, much subsequent work has centered around the narrow passage problem [3, 28]. Due to the intractable nature of motion planning, tangible improvements often came from making assumptions about the underlying space, such that the human engineer could encode case specific heuristics into the search. Some work, rather than improve runtime, focused attention on the types of solutions output by the algorithm, for example searching for paths with high clearance [16].
@@ I Introduction
In formally discussing guiding spaces we make the source of guidance explicit. Doing so allows us to identify guidance-generating components in many seemingly distinct prior methods, which fit under three main subcategories; robot modification, environment modification, and experience based guidance. The perspective of guiding spaces helps isolate conceptual contributions of different algorithms, and consequently yields simpler generalized implementations that are also easily composed. Framing SBMP algorithms in this common language enables us to define a new method for evaluation that focuses on the quality of guidance.
@@
We approach the task of evaluating guidance from an information theoretic perspective, whereby the procedure for computing guidance conveys information to the search process. We test our metric experimentally on a range of algorithms and environments. The results match intuition for the suitability of each algorithm to features in certain c-spaces. Moreover, the metric captures properties of algorithms otherwise obscured by traditional metrics such as runtime or number of samples. Measuring quality of guidance locally gives insight into the value of combining guiding spaces, which we demonstrate with a hybrid guiding space that outperforms its individual parts.
II Preliminaries
@@
A motion planning problem consists of the tuple $M=(r,E,s,t)$, for robot $r$, environment $E$, and task $(s,t)$, where the objective is to find the shortest collision free path $P_{s,t}$ from starting configuration $s$ to goal configuration $t$. Note that $M$ is the single-query variant of the motion planning problem, and we will touch on the multi-query version later in this section. Together, $(r,E)$ describe the configuration space $\mathcal{C}$, a continuous space consisting of all possible configurations of $r$ in $E$. This space is implicitly partitioned into two parts, $C_{free}$ and $C_{obs}$. $C_{free}$ is the set of all configurations in which $r$ does not intersect with either itself or obstacles in $E$, and $C_{obs}=\mathcal{C}\setminus\mathcal{C}_{free}$.
@@
SBMP methods, such as RRT, explore $\mathcal{C}$ to solve $M$ by building a search tree $T$. They do so by iteratively picking an existing node $v\in T$, and then expanding the tree by adding a new node as a child of $v$. We refer to these two operations as node selection, which is performed by a sampler, and node expansion, which is performed by a local planner. We refer to the set of points reachable by the local planner as $N_{\mathcal{C}}(v)$ for $v\in T$, the neighborhood of $v$. Usually the initial tree contains only the start node $s$ and the algorithm terminates when a configuration sufficiently close to $t$ is added to the tree.
@@
Note that in this work we depart from the classical partition into tree and roadmap algorithms (e.g., PRM [22]). Our perspective is that roadmap based algorithms (which build a general graph structure instead of a search tree) preprocess $\mathcal{C}$ to produce a data structure that solves the multi-query version of the problem. In particular, most variants of roadmap algorithms do not depend on the task $(s,t)$. In other words, roadmap algorithms do not solve single task search problems such as $M$ but rather provide guidance for doing so in a given c-space, where the roadmap is the guiding space. This perspective motivates our definition of guided motion planning: a tree (or forest) based search algorithm that uses a guiding space to inform exploration.
III Defining Guided Planning
(a) A 2D workspace with a stationary 2-joint manipulator, and 2 obstacles. Three intermediates from the path found in the guiding space (b) are shown in dashed gray, where the last is the target configuration. We observe that one intermediate is invalid.
(b) The lazy planning guiding space for the environment (a) without obstacle $o_{2}$. A red curve shows a path between the current and target configurations shown in (a), and a blue wedge shows the resulted sampling bias.
Figure 1: An illustration of guiding space usage.
@@ III Defining Guided Planning
We define a search space $\mathcal{X}$ as a state space (set) with a distance function $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$. In contrast, we define a guiding space $\mathcal{S}$ as a state space which assigns a value to each state, namely there exists some function $h_{\mathcal{S}}:\mathcal{S}\rightarrow\mathbb{R}_{+}$, which we call the heuristic. Note that a c-space $\mathcal{C}$ is a search space and that every search space can be viewed as a guiding space, i.e., $h_{t}(x)=d(x,t)$ for $x\in\mathcal{C}$. In other words, distance to some state (goal) $t$ can be used as an heuristic.
Consider the motion planning problem $M=(r,E,s,t)$ depicted in Figure 1, where a two joint planar manipulator $r$ must navigate from state $s$ to $t$ while avoiding obstacles $E=\{o_{1},o_{2}\}$. Now consider the related c-space in Figure 1(b), $\mathcal{C}^{\prime}=(r,E^{\prime}),E^{\prime}=\{o_{1}\}$. Let $\mathcal{S}$ be the guiding space whose heuristic is the distance to $t$ in $\mathcal{C}^{\prime}$, i.e., $h_{\mathcal{S}}(x)=d_{\mathcal{C}^{\prime}}(x,t)$. As we will show next, this guiding space can be used to guide planning in $\mathcal{C}$.
@@
A guiding space projection $P_{T}:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{S}$ is a mapping from pairs of states in some search space $\mathcal{X}$ to states of a guiding space $\mathcal{S}$. This mapping is conditioned on a search tree $T$, which intuitively represents the known information regarding $\mathcal{C}_{free}$ and $\mathcal{C}_{obs}$. By relating tasks in $\mathcal{X}$ to states in $\mathcal{S}$ the projection allows a guided search algorithm to use the heuristic during search. Often for convenience we will drop the second term to the projection, namely $P_{T}(x)=P_{T}(x,t)$.
In the example of Figure 1 the projection is $P_{T}((x,t)\in\mathcal{C}\times\mathcal{C})=x\in\mathcal{C}^{\prime}$. Note that this projection does not change the actual configuration $x$, and only modifies its validity since configurations which collide only with obstacle $o_{2}$ in $\mathcal{C}$ are valid in $\mathcal{C}^{\prime}$. Moreover, notice that despite colliding with obstacle $o_{2}$, the path highlighted in red still contains useful information for computing paths in $\mathcal{C}$. First, the length of this path is a good estimate for the true global shortest path length, since a slight deviation is collision free and optimal. Second, regardless of global properties of the path, at certain configurations it contains locally near-optimal controls. The key insight is that not only does this path contain useful information, but also it is easier to compute than the true optimal solution, since it corresponds to a simpler search problem. The existence of quick to compute useful information is what motivates our definition of guided motion planning.
@@
A guided motion planning problem $M_{P}=(r,E,s,t,P)$ is a motion planning problem augmented with a projection $P$. We note that the image of $P$ describes a (potentially implicit) guiding space $\mathcal{S}$ and corresponding $h$. Next we formalize both how this information is used and where it comes from.
@@ III-A Guided Search
Given a search tree $T$, a baseline exploration algorithm might explore the space by selecting an existing state on the tree uniformly at random and then expanding from it according to some local planner. Unfortunately this approach has poor exploration properties, as it is heavily biased towards regions that have already been explored. Discrete search avoids this problem as newly explored states are checked against a cached set of visited states.
@@
Algorithm 1 shows a guided search procedure for solving a problem $M_{P}$. In each iteration the algorithm maps nodes on the tree through the projection $P_{T}$ then uses $h$ to determine the next node to explore, i.e., by selecting the node with minimum $h$. In general, we can think of the projection as providing guidance by computing a distribution over the nodes of the tree from which we can sample to select how to expand. The tree then grows according to the neighborhood defined by the local planner. This parallels the well known heuristic guided discrete search algorithm A*.
III-B Computing a Guiding Space Projection $P$
Consider an unknown but fixed distribution $\mathcal{D}$ over robots, environments, and tasks. Guided search consists of an initial offline training phase, during which we train a projection $P^{M}_{T}$ on samples $\{M~{}|~{}M\sim\mathcal{D}\}$ from this distribution. Note we are generally not concerned with time complexity during training, and that many algorithms skip this phase entirely. For example, RRT does not “train”, and performs the same procedure regardless of c-space. More commonly, after making assumptions about the c-space distribution, the human engineer encodes their prior knowledge into a suitable procedure for computing the projection, thus replacing the training phase. For example knowing that most obstacles in the environment are fixed might motivate projecting onto a guiding space in which collision with those obstacles has been precomputed [21]. Alternatively, knowing that some degree of freedom has negligible impact on $\mathcal{C}$ might motivate projecting onto the other dimensions [24].
Next begins the online phase, in which a specific instantiation $M\sim\mathcal{D}$ of a motion planning problem is sampled from the distribution and we conduct guided search on $M_{P_{T}}$ following Algorithm 1.
III-C Simple Examples of Guiding Spaces
@@
One guiding space for $\mathcal{C}$ is itself, $\mathcal{S}=\mathcal{C}$, where $P_{T}(x)=x$ is the identity and $h=d_{\mathcal{C}}(x,t)$ is the true distance to the goal configuration. This highlights the general setting in which a search space is used as a guiding space, in which for guidance to be useful we expect that solving tasks in $\mathcal{S}$ be easier than in $\mathcal{C}$. Similarly, $\mathcal{S}=\{0\}$ is a guiding space with a trivial projection $P$ and uninformed $h$. Such trivial projections correspond to fixed sampling distributions over $T$. We observe a natural trade-off between guidance quality and computation difficulty - computing solutions when $\mathcal{S}=\mathcal{C}$ is difficult (the original problem) but they provide perfect guidance, whereas when $\mathcal{S}=\{0\}$ we can compute $h$ in constant time but get uninformed guidance.
@@
The classic implementation of RRT samples uniformly in $\mathcal{C}$ to select which node on the tree to expand. This is clearly not equivalent to sampling a node uniformly at random from the tree, and is known to consequently guide exploration towards unexplored regions via Voronoi bias. While in practice the guiding space is implicit in the sampling and nearest neighbor procedure, we can compute it explicitly through the Voronoi decomposition, where the value $h$ of each node on the tree is proportional to the volume of its corresponding cell. Notice how the guiding space depends on the search tree, as the nodes dictate the Voronoi cells.
@@
As noted previously, a PRM in configuration space is a guiding space. If the distribution $\mathcal{D}$ contains a single robot and environment, but is stochastic over tasks, then precomputing a discrete roadmap in $\mathcal{C}$ is a common approach for answering multiple queries. The projection $P$ is often taken to be the function which returns the nearest neighbor on the graph. For any task $(s,t)$ and vertex $v=P(s)$, the heuristic $h_{\mathcal{S}}(v)$ is the length of the shortest path on the graph between $v$ and $P(t)$. This length is an approximation to the optimal solution for that task and can sometimes provide imperfect guidance; for example not every point in space might be locally navigable to a node on the graph.
IV Related work: categories of guidance
In this section we discuss existing work which does guided search. Note that many of the works cited below do not discuss guidance as a contribution, yet here we identify that they do implicitly perform guided search.

### v2

@@ Abstract
1 Introduction
A general paradigm in artificial intelligence is that one can solve provably worst case intractable search problems using heuristics that are effective in the average case, i.e., in most real world problems. This same principle holds true in robotics, where much of the difficulty stems from the geometry and the implicit nature of the search space, called the configuration space (C-space) [31]. The family of sampling based motion planning (SBMP) algorithms searches the continuous C-space using random sampling and local extensions in order to find valid paths.
Most such algorithms implicitly perform guided search. Early algorithms, such as Rapidly-exploring Random Trees (RRT) [30], showed that exploration through Voronoi bias, i.e., choosing which node to expand with probability proportional to the volume of its Voronoi cell, is highly effective in the absence of arbitrarily narrow passages [27, 24]. Much work has centered around the narrow passage problem [3, 38]. Due to the intractable nature of motion planning [37], tangible improvements often came from making assumptions about the underlying space, such that the human engineer could encode case specific heuristics to bias the search. Some work, rather than improving runtime, focused attention on the types of solutions output by the algorithm, for example searching for paths with high clearance [18, 41, 43, 20, 22].
@@ 1 Introduction
In formally discussing guiding spaces we make the source of guidance explicit. Doing so allows us to identify guidance-generating components in many seemingly distinct prior methods, which fit under three main subcategories; robot modification, environment modification, and experience-based guidance. The perspective of guiding spaces helps isolate conceptual contributions of different algorithms and consequently yields simpler generalized implementations. These new implementations are more easily composed into algorithms that take advantage of the strengths of independent methods. Another benefit of framing SBMP algorithms in this common language is that it enables us to define a new method for evaluation that focuses on the quality of guidance.
@@
We approach the task of evaluating guidance from an information theoretic perspective, whereby the procedure for computing guidance conveys information to the search process. We test our metric experimentally on a range of algorithms and environments. The results match intuition for the suitability of each method to features in certain C-spaces. Moreover, the metric captures properties otherwise obscured by the traditional metrics of runtime or number of samples, such as sensitivity to local environment features. Measuring quality of guidance locally gives insight into the value of combining guiding spaces, which we demonstrate with a hybrid guiding space that outperforms its individual parts.
2 Preliminaries
@@
A motion planning problem consists of the tuple $M=(r,E,s,t)$, for robot $r$, environment $E$, and task $(s,t)$, where the objective is to find a (possibly shortest) valid path $P_{s,t}$ from starting configuration $s$ to goal configuration $t$. Note that $M$ is the single-query variant of the motion planning problem, and we will touch on the multi-query version later in this section. Together, $(r,E)$ describe the configuration space $\mathcal{C}$, a continuous space consisting of all possible configurations of $r$ in $E$. This space is implicitly partitioned into two parts, $C_{free}$ and $C_{obs}$. $C_{free}$ is the set of all valid configurations, and $C_{obs}=\mathcal{C}\setminus\mathcal{C}_{free}$.
@@
Incremental SBMP methods, such as RRT, explore $\mathcal{C}$ to solve $M$ by building a search tree $T$. They do so by iteratively picking an existing node $v\in T$, and then expanding the tree by adding a new node as a child of $v$. We refer to these two operations as node selection, which is performed by a sampler, and node expansion, which is performed by a local planner. We refer to the set of points reachable by the local planner as $N_{\mathcal{C}}(v)$ for $v\in T$, the neighborhood of $v$. Usually the initial tree contains only the start node $s$ and the algorithm terminates when a configuration sufficiently close to $t$ is added to the tree.
@@
Note that in this work we depart from the classical partition into tree and graph-based (roadmap) algorithms (e.g., PRM [28]). Most variants of graph-based algorithms attempt to solve the multi-query version of the problem. Our perspective is that such methods take advantage of a powerful prior; that future motion planning instances will be in the same C-space $\mathcal{C}$ but have different tasks. As such, preprocessing $\mathcal{C}$ to produce a data structure (the graph) which guides future motion planning queries by reducing them to path planning is highly effective. This perspective motivates our definition of guided motion planning: a tree based search algorithm that uses prior knowledge to generate a guiding space which informs exploration.
3 Defining Guided Planning
Figure 1: An illustration of guiding spaces and their usage. (a) depicts a 2D workspace with a fixed-base 2-joint manipulator and 2 obstacles, alongside two configurations $s$ and $t$ and two intermediates (in dashed gray) on the $(s,t)$-path found in the guiding space (b). We observe that one intermediate is invalid. In (b) we see the lazy planning guiding space for the environment (a) without obstacle $o_{2}$. A red curve shows a path between the current and target configurations shown in (a), and a blue wedge shows the resulting sampling bias. In (c) we see another 2D environment with a 3 DoF $(x,y,\theta)$ rectangular robot, and in (d) we see a guiding space, the workspace medial-axis, and the projections of the configurations $s$ and $t$ shown in (c) onto the medial-axis. Again, note that the shortest $(s,t)$ path in (d) is not valid in (c).
In this section, we define the main components of guided motion planning: namely the guiding space, a mapping which relates C-space to the guiding space, and an algorithm that uses both to plan. We provide some simple intuitive examples followed by a categorization of guiding spaces in the next section.
At a high level we think of guidance as the knowledge distilled from past experience to be used in future motion planning. Thus, in the setting of guided motion planning, we assume there exists an unknown but fixed distribution over robots, environments, and tasks. Guided search consists of an optional offline training phase, during which we learn from this distribution, followed by an online phase, in which a specific motion planning problem is sampled from the distribution and solved. We will also show that a search space can be used as a guiding space, making the definition naturally hierarchical since that search space can, in turn, have guidance.
Note, we are generally not concerned with time complexity during the offline phase, and that many algorithms skip it entirely. For example, RRT does not “learn”, and performs the same procedure regardless of C-space. More commonly, after making assumptions about the C-space distribution, the human engineer encodes their prior knowledge into a suitable procedure for computing guidance, thus replacing the training phase. For example, knowing that some obstacles in the environment are fixed might motivate precomputing collision with those obstacles [26] (where PRM is an extreme version of this). Alternatively, knowing that some degree of freedom has negligible impact on $\mathcal{C}$ might motivate planning with only the other dimensions [32, 42].
We now define the individual components of guided motion planning to make explicit where and how guidance is used.
3.1 Guided Planning Definition
@@
We define a search space $\mathcal{X}$ as a state space (set) with a distance function $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$. In contrast, we define a guiding space $\mathcal{S}$ as a state space which assigns a value to each state, namely there exists some function $h_{\mathcal{S}}:\mathcal{S}\rightarrow\mathbb{R}_{+}$, which we call the heuristic. Note that a C-space $\mathcal{C}$ is a search space and that every search space can be viewed as a guiding space, i.e., $h_{t}(x)=d(x,t)$ for $x\in\mathcal{C}$. In other words, distance to some goal state $t$ can be used as a heuristic.
@@
A guiding space projection $P_{T}:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{S}$ is a mapping from pairs of states (tasks) in some search space $\mathcal{X}$ to states of a guiding space $\mathcal{S}$. This mapping is conditioned on a search tree $T$, which intuitively represents the known information regarding $\mathcal{C}_{free}$ and $\mathcal{C}_{obs}$. By relating tasks in $\mathcal{X}$ to states in $\mathcal{S}$ the projection allows a guided search algorithm to use the heuristic during search. Often for convenience we will drop the second term to the projection, namely $P_{T}(x)=P_{T}(x,t)$. The offline training phase of guided motion planning produces a procedure $P^{M}_{T}$ for computing a projection for a new motion planning problem by training on samples $M\sim\mathcal{D}$ for some distribution $\mathcal{D}$ over robots, environments, and tasks.
@@
A guided motion planning problem $M_{P}=(r,E,s,t,P)$ is a motion planning problem augmented with a projection $P$. We note that the image of $P$ describes a (potentially implicit) guiding space $\mathcal{S}$ and corresponding $h$.
Consider the motion planning problem $M=(r,E,s,t)$ depicted in Figure 1(a), where a two joint planar manipulator $r$ must navigate from state $s$ to $t$ while avoiding obstacles $E=\{o_{1},o_{2}\}$. Now consider the related C-space in Figure 1(b) which omits obstacle $o_{2}$, $\mathcal{C}^{\prime}=(r,E^{\prime}),E^{\prime}=\{o_{1}\}$. Let $\mathcal{S}$ be the guiding space whose heuristic is the distance to $t$ in $\mathcal{C}^{\prime}$, i.e., $h_{\mathcal{S}}(x)=d_{\mathcal{C}^{\prime}}(x,t)$. In this example the guiding space projection is $P_{T}((x,t)\in\mathcal{C}\times\mathcal{C})=x\in\mathcal{S}_{t}$, where $h_{\mathcal{S}_{t}}(x)=d_{\mathcal{C}^{\prime}}(x,t)$. Notice that it does not depend on the search tree, and uses distances in $\mathcal{C}^{\prime}$ as heuristics (to guide search in $\mathcal{C}$). In addition, while the projection may seem like the identity mapping $x\mapsto x$, it is implicitly modifying the validity of $x$ since configurations which collide only with obstacle $o_{2}$ in $\mathcal{C}$ are valid in $\mathcal{C}^{\prime}$. Despite colliding with obstacle $o_{2}$, the path highlighted in red still contains useful information for computing paths in $\mathcal{C}$. Not only is the length of this path a good estimate for the true shortest path length, but also at certain configurations it contains locally near-optimal controls. The key insight for guidance is that not only does this path contain useful information, but also it is easier to compute than the true optimal solution, since it corresponds to a simpler search problem.
@@ 3.2 Guided Search
Given a search tree $T$, a baseline exploration algorithm might explore the space by selecting an existing state on the tree uniformly at random and then expanding from it according to some local planner. Unfortunately this approach has poor exploration properties[30], as it is heavily biased towards regions that have already been explored. Discrete search avoids this problem as newly explored states are checked against a cached set of visited states.
@@
Algorithm 1 shows an example procedure for the online phase of guided search for solving a problem $M_{P}$. In each iteration the algorithm maps nodes on the tree through the projection $P_{T}$ then uses $h$ to determine the next node to explore, i.e., by selecting the node with minimum $h$. In general, we can think of the projection as providing guidance by computing a distribution over the nodes of the tree from which we can sample to select how to expand. The tree then grows according to the neighborhood defined by the local planner.
This parallels the well known heuristic guided discrete search algorithm A*, where the main difference is that instead of maintaining an explicit ordered frontier for search, we allow the guiding space to update its distribution over $T$ at each iteration. This distinction is crucial - each new sample provides information regarding which regions of space to explore such that no fixed ordering of states (e.g., A*) is an effective exploration strategy.
3.3 Simple Examples of Guiding Spaces
@@
One guiding space for $\mathcal{C}$ is itself, $\mathcal{S}=\mathcal{C}$, where $P_{T}(x)=x$ is the identity and $h=d_{\mathcal{C}}(x,t)$ is the true distance to the goal configuration. This highlights the general setting in which a search space is used as a guiding space, in which, for guidance to be useful, we expect that solving tasks in $\mathcal{S}$ should be easier than in $\mathcal{C}$. Similarly, $\mathcal{S}=\{0\}$ is a guiding space with a trivial projection $P$ and uninformed $h$. Such trivial projections correspond to fixed sampling distributions over $T$.
We observe a natural trade-off between guidance quality and computation difficulty - computing solutions when $\mathcal{S}=\mathcal{C}$ is difficult (the original problem) but they provide perfect guidance, whereas when $\mathcal{S}=\{0\}$ we can compute $h$ in constant time but get uninformed guidance.
@@
The classic implementation of RRT samples uniformly in $\mathcal{C}$ to select which node on the tree to expand. This is not equivalent to sampling a node uniformly at random from the tree, and is known to guide exploration towards unexplored regions via Voronoi bias. While in practice the guiding space is implicit in the sampling and nearest neighbor procedure, we can compute it explicitly through the Voronoi decomposition, where the value $h$ of each node on the tree is (inversely) proportional to the volume of its corresponding cell. Notice how the guiding space depends on the search tree, as the nodes dictate the Voronoi cells.
@@
As mentioned previously, a PRM in configuration space is a guiding space. If the distribution $\mathcal{D}$ contains a single robot and environment, but is stochastic over tasks, then precomputing a discrete roadmap in $\mathcal{C}$ is a common approach for answering multiple queries. The projection $P$ is often taken to be the function which returns the nearest neighbor on the graph. For any task $(s,t)$ and vertex $v=P(s)$, the heuristic $h_{\mathcal{S}}(v)$ is the length of the shortest path on the graph between $v$ and $P(t)$. This length is an approximation to the optimal solution for that task and can sometimes provide imperfect guidance; for example not every point in space might be locally navigable to a node on the graph.
One of the most commonly used forms of guidance comes from the workspace (see Figure 1(d)). Planning for a point robot in the low-dimensional workspace is relatively easy, and while not every workspace path is feasible in C-space, every C-space path corresponds to some feasible path in workspace. In other words, in the context of A*, we can think of distance in workspace as an admissible heuristic for distance in C-space (with some possible scaling). In the next section we describe how workspace guidance methods can be thought of as using plans for a modified version of the robot to guide planning in the original C-space.
4 Categories of Guidance